# GeoSplat: A Deep Dive into Geometry-Constrained Gaussian Splatting

### v2

@@ Abstract
A few recent works explored incorporating geometric priors to regularize the optimization of Gaussian splatting, further improving its performance. However, those early studies mainly focused on the use of low-order geometric priors (e.g., normal vector), and they are also unreliably estimated by noise-sensitive methods, like local principal component analysis. To address their limitations, we first present GeoSplat, a general geometry-constrained optimization framework that exploits both first-order and second-order geometric quantities to improve the entire training pipeline of Gaussian splatting, including Gaussian initialization, gradient update, and densification. As an example, we initialize the scales of 3D Gaussian primitives in terms of principal curvatures, leading to a better coverage of the object surface than random initialization. Secondly, based on certain geometric structures (e.g., local manifold), we introduce efficient and noise-robust estimation methods that provide dynamic geometric priors for our framework. We conduct extensive experiments on multiple datasets for novel view synthesis, showing that our framework: GeoSplat, significantly improves the performance of Gaussian splatting and outperforms previous baselines.
@@ 1 Introduction
The photorealistic rendering quality and high efficiency of Gaussian splatting (Kerbl et al., 2023) has sparked a surge of recent studies, which further it from various perspectives, such as memory consumption (Niedermayr et al., 2024) and texture (Gao et al., 2024).
@@
A notable perspective from those studies is to treat geometric priors as certain regularization for improving Gaussian splatting, given that the Gaussian primitives should approximate the surfaces of 3D objects. For instance, Li et al. (2025) placed new primitives in the tangent plane of the original one, with the aim to reduce outlier artifacts. While obtaining performance gains, previous methods mainly focused on low-order geometric quantities (e.g., normal vector) and did not much consider higher-order ones (i.e., curvature). The two types of geometric information characterize very different aspects of 2D object surfaces (Lee, 2018), so the absence of either one might limit their potential.
Another key limitation of prior works is that they estimated geometric information through unreliable methods. For example, local principal component analysis (PCA) (Kambhatla & Leen, 1997; Li et al., 2025) is highly sensitive to noise, and the trained StableNormal model (Ye et al., 2024; Wang et al., 2024) often fails on rarely seen data. The geometric priors in previous methods were also static, which would gradually become inaccurate in characterizing the dynamic geometry of Gaussian primitives during optimization.
@@
To address the limitations of prior works, we first propose a general geometry-constrained optimization framework: GeoSplat, exploiting both low-order and high-order geometric priors to regularize the training pipeline of Gaussian splatting. While being largely ignored by previous methods, the high-order information (e.g., curvature) characterizes a critical property of 2D surface: how it bends in the 3D space. In light of this, we adopt such information to regularize the shapes of Gaussian primitives. For instance, an area with low curvature indicates that it approximates a flat plane, so we can initialize the scales of Gaussian primitives in this area as a large number. Low-order geometric quantities, such as tangent and normal vectors, have already played key roles in recent works, and we further extend its application scopes in our framework. For example, to reduce floating artifacts, we truncate the gradient update of a Gaussian primitive in terms of its normal direction.
Secondly, we present an estimation method that can provide noise-robust geometric information for our framework. Specifically, we assume each Gaussian primitive resides in an underlying surface that is locally a manifold (Lee, 2018), and derive the analytical form of shape operator that contains the curvature information. We also explore another language in geometry: varifold (Buet et al., 2017), which outperforms our manifold-based method at times. Both methods are highly efficient, and thus can provide our framework with dynamic geometric priors during training.
Extensive experiments have been conducted on multiple benchmark datasets in novel view synthesis. The results demonstrate that our framework: GeoSplat, significantly improves the performance of Gaussian splatting and outperforms previous baselines.
@@ 2.1 Gaussian Splatting
The core of Gaussian splatting is to represent a 3D scene as a number of Gaussian primitives $\{\mathcal{G}^{i}_{\mathrm{obj}}\}_{i\in\mathcal{I}}$ indexed by set $\mathcal{I}$, with each parameterized by its opacity $\alpha^{i}_{\mathrm{obj}}\in\mathbb{R}$, RGB color $\mathbf{c}^{i}_{\mathrm{obj}}\in\mathbb{R}^{3}$, mean vector $\bm{\mu}^{i}_{\mathrm{obj}}\in\mathbb{R}^{3}$, and covariance matrix $\bm{\Sigma}^{i}_{\mathrm{obj}}\in\mathbb{R}^{3\times 3}$. The matrix $\bm{\Sigma}^{i}_{\mathrm{obj}}$ is semi-positive definite and has a decomposed form: $\mathbf{R}^{i}_{\mathrm{obj}}\mathbf{S}^{i}_{\mathrm{obj}}(\mathbf{R}^{i}_{\mathrm{obj}}\mathbf{S}^{i}_{\mathrm{obj}})^{\top}$, where $\mathbf{S}^{i}_{\mathrm{obj}}=\mathrm{diag}([s^{i}_{1},s^{i}_{2},s^{i}_{3}]^{\top})$ is the diagonal scale matrix and $\mathbf{R}^{i}_{\mathrm{obj}}=[\mathbf{r}_{1}^{i};\mathbf{r}_{2}^{i};\mathbf{r}_{3}^{i}]\in\mathrm{SO}(3)$ is the rotation matrix.
@@
where $\mathbf{T}_{\mathrm{pose}}\in\mathrm{SE}(3)$ is the camera pose consisting of rotation matrix $\mathbf{W}_{\mathrm{pose}}$ and translation vector $\mathbf{b}_{\mathrm{pose}}$, and $\Pi$ is the projection map from a 3D coordinate to 2D one, with $\mathbf{J}$ being its Jacobian matrix. Then, the blending operation renders the color of pixel $\mathbf{p}$ as
@@
where $\beta^{k}(\mathbf{p})$ quantifies the radiance of a Gaussian primitive: $\alpha^{k}_{\mathrm{obj}}\mathcal{N}(\mathbf{p};\bm{\mu}^{k}_{\mathrm{plane}},\bm{\Sigma}^{k}_{\mathrm{plane}})$. Under this scheme, the optimization objective is to match rendered with real images.
@@ 2.2 Problem Formulation
Our primary goal is to incorporate diverse geometric priors into optimizing Gaussian splatting, especially the high-order ones (e.g., curvature). Meanwhile, we also aim to estimate the priors in a noise-robust manner. These two aspects are largely neglected in recent explorations (Wang et al., 2024; Li et al., 2025) on geometric regularization.
@@
Gaussian primitives $\{\mathcal{G}^{i}_{\mathrm{obj}}\}_{i\in\mathcal{I}}$ reside in an underlying 3D scene formed by finitely many continuous manifold $\{\mathcal{M}_{j}\}_{j\in\mathcal{J}}$, with index set $\mathcal{J}$. Notably, each manifold $\mathcal{M}_{j}$ is smooth up to a zero-measure set of singular points.
@@
Compared with vanilla manifold, the above new setting can accommodate typical 3D scenes, where there are multiple independent spatial objects that might not be fully smooth. The zero-measure condition is also very important, permitting almost every point $\mathbf{q}\in\mathbb{R}^{3}$ in the scene to locate at a local region that can be treated as a smooth manifold. For novel view synthesis, the surface $\mathcal{M}_{j}$ is normally 2D: $D_{\mathcal{M}_{j}}=2$, embedded in the 3D ambient space: $\mathcal{E}=\mathbb{R}^{3}$.
@@
Most geometric quantities are locally defined, and thus fit our setting. Specifically, every Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}},i\in\mathcal{I}$ will lie on a certain underlying object surface $\mathcal{M}_{j},j\in\mathcal{J}$ at position $\mathbf{q}=\bm{\mu}^{i}_{\mathrm{obj}}$, with tangent space $\mathcal{T}_{\mathbf{q}}\mathcal{M}_{j}$ spanned by orthonormal basis $\{\mathbf{u}^{i}_{d}\}_{1\leq d\leq D_{\mathcal{M}_{j}}}$ and normal space $\mathcal{N}_{\mathbf{q}}\mathcal{M}_{j}$ formed by another basis $\{\widetilde{\mathbf{u}}^{i}_{d}\}_{D_{\mathcal{M}_{j}}+1\leq d\leq D_{\mathcal{E}}}$. These vector bases belong to first-order geometric quantities, and notably, we also aim to consider high-order ones: principal curvatures and directions $\{(\tau_{d}^{i},\widetilde{\mathbf{w}}^{i}_{d})\}_{1\leq d\leq D_{\mathcal{M}_{j}}}$, with each pair indicating that the underlying surface is bent with a magnitude of $\tau_{d}^{i}$ along direction $\widetilde{\mathbf{w}}^{i}_{d}$. For notational convenience, we will omit the subscript $j$ of local manifold $M_{j}$, since it is unique for each primitive $\mathcal{G}^{i}_{\mathrm{obj}}$.
Another geometric structure: varifold.
@@
We also explore adopting another language in geometry: varifold (Allard, 1972), which is free from the smoothness assumption of manifold. Simply speaking, it is defined as a joint non-negative Radon measure (Rudin, 1987) $\mathcal{W}=\theta\mathcal{H}^{D_{\mathsf{M}}}\otimes\delta_{\mathcal{T}_{\mathbf{\star}}\mathsf{M}}$ on the product space of a rectifiable set (Federer, 2014) $\mathsf{M}$ and the Grassmannian (Lee, 2018) $\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$. Here $\theta:\mathsf{M}\rightarrow\mathbb{R}$ is a non-negative function, $\delta_{\mathcal{T}_{\mathbf{\star}}\mathsf{M}}$ represents a type of Dirac measure, and $\mathcal{H}^{D_{\mathsf{M}}}$ denotes the Hausdorff measure (Federer, 2014).
@@
Varifold is so flexibly defined that we can directly model the discrete locations of Gaussian primitives $\{\mathcal{G}^{i}_{\mathrm{obj}}\}_{i\in\mathcal{I}}$ as the set $\mathsf{M}$, without supposing there is any smooth interpolation. More importantly, the geometric quantities defined for manifold can be generalized for varifold, such as the approximate tangent space $\mathcal{T}_{\mathbf{\star}}\mathsf{M}$ and generalized mean curvature $\mathsf{H}$. Due to the limited space of main text, we provide their details, with a concise review of varifold, in Appendix A.2.
@@ 3 Method: Geometry-constrained Gaussian Splatting
In this section, we first present GeoSplat, a framework of geometry-constrained Gaussian splatting, which exploits high-order geometric priors that are largely neglected by previous works. Then, we present noise-robust methods to estimate dynamic geometric priors. In contrast, previous methods obtained static priors through noise-sensitive or domain-specific techniques.
@@ 3.1.1 Curvature-guided Gaussian Initialization
We assume that the primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ locates at an underlying surface $\mathcal{M}$ [^1] With abuse of notation, we mostly denote the 2D surface as $\mathcal{M}$, potentially assuming it is a local manifold. However, it should be denoted as a rectifiable set $\mathsf{M}$ in the context of varifold., so it is intuitive to warm up its initial shape $\mathbf{\Sigma}^{i}_{\mathrm{obj}}$ in terms of geometric priors. To begin with, we set one rotation direction as the normal vector, with the corresponding scale specified as a tiny value:
$\mathbf{r}_{3}^{i}=\widetilde{\mathbf{u}}^{i}_{3},\ \ \ s^{i}_{3}=\xi_{\mathrm{min}}\approx 0.$ (3)
@@
If MAC is quite small: $\widebar{\tau}^{i}<\xi_{\mathrm{min}}$, which means the curvature is basically negligible, then we can identify that the corresponding Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ resides in an almost flat region. In this case, we upsample its position through midpoint interpolation:
@@ 3.1.2 Shape-constrained Optimization
where $\omega$ is the learning rate, $\nabla^{i}\mathfrak{L}$ denotes the gradient of loss function $\mathfrak{L}$ regarding position $\bm{\mu}^{i}_{\mathrm{obj}}$, $(\cdot)^{\top}$ means to project a vector to the tangent plane $\mathcal{T}_{\mathbf{q}}\mathcal{M}$, and $(\cdot)^{\perp}$ represents the projection to normal direction $\widetilde{\mathbf{u}}^{i}_{3}$. With this truncation, a Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ that lies on the underlying surface $\mathcal{M}$ will still stay close to it after an aggressive gradient update $\nabla^{i}\mathfrak{L}$.
@@
Another type of artifact that might degrade the performance of Gaussian splatting is the Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ with a needle-like shape (Hyung et al., 2024): covariance matrix $\mathbf{\Sigma}^{i}_{\mathrm{obj}}$ degenerates to be rank-$1$. From a geometric perspective, the main scales $s^{i}_{1},s^{i}_{2}$ of the primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ needs to be regularized to ensure a rank-$2$ covariance matrix $\mathbf{\Sigma}^{i}_{\mathrm{obj}}$, so that it will have a sufficient coverage over the underlying manifold $\mathcal{M}$.
@@
with the last term to ensure the rank of matrix $\mathbf{\Sigma}^{i}_{\mathrm{obj}}$ is smaller than $3$: the Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ is closely attached to the surface $\mathcal{M}$. For the hinge term, we again regard the curvature information $\{\tau_{d}^{1},\tau_{d}^{2}\}$ as natural guidance for shaping the scale matrix $\mathbf{S}^{i}_{\mathrm{obj}}$. In a similar spirit, we introduce another regularization for the rotation matrix $\mathbf{R}^{i}_{\mathrm{obj}}$ as

### v3

@@ Abstract
A few recent works explored incorporating geometric priors to regularize the optimization of Gaussian splatting, further improving its performance. However, those early studies mainly focused on the use of low-order geometric priors (e.g., normal vector), and they might also be unreliably estimated by noise-sensitive methods, like local principal component analysis. To address their limitations, we first present GeoSplat, a general geometry-constrained optimization framework that exploits both first-order and second-order geometric quantities to improve the entire training pipeline of Gaussian splatting, including Gaussian initialization, gradient update, and densification. As an example, we initialize the scales of 3D Gaussian primitives in terms of principal curvatures, leading to a better coverage of the object surface than random initialization. Secondly, based on certain geometric structures (e.g., local manifold), we introduce efficient and noise-robust estimation methods that provide dynamic geometric priors for our framework. We conduct extensive experiments on multiple datasets for novel view synthesis, showing that our framework, GeoSplat, significantly improves the performance of Gaussian splatting and outperforms previous baselines.
@@ 1 Introduction
The photorealistic rendering quality and high efficiency of Gaussian splatting (Kerbl et al., 2023) have fueled a surge of recent studies, which further investigate it from various perspectives, such as memory consumption (Niedermayr et al., 2024) and texture (Gao et al., 2024).
@@
A notable perspective from those studies is to treat geometric priors as a form of regularization for improving Gaussian splatting, given that the Gaussian primitives should approximate the surfaces of 3D objects. For example, Li et al. (2025) places new primitives in the tangent planes of original primitives, with the aim to reduce outlier artifacts. While obtaining performance gains, previous methods mainly focus on low-order geometric quantities (e.g., normal vector) and do not much consider higher-order ones (i.e., curvature). The two types of geometric information characterize very different aspects of 2D object surfaces (Lee, 2018), so the absence of either one might limit their potential.
Another key limitation of prior works is that they sometimes estimate geometric information using methods that lack robustness. For example, local principal component analysis (PCA) (Kambhatla & Leen, 1997; Li et al., 2025) is highly sensitive to noise, and the trained StableNormal model (Ye et al., 2024; Wang et al., 2024) often fails on rarely seen data. The geometric priors in previous methods are also static, which would gradually become inaccurate in characterizing the dynamic geometry of Gaussian primitives during optimization.
@@
To address the limitations of prior works, we first propose a general geometry-constrained optimization framework, GeoSplat, exploiting both low-order and higher-order geometric priors to regularize the training pipeline of Gaussian splatting. While being largely ignored by previous methods, the higher-order information (e.g., curvature) characterizes a critical property of 2D surface: how it bends in the 3D space. In light of this, we adopt such information to regularize the shapes of Gaussian primitives. For instance, an area with low curvature indicates that it approximates a flat plane, so we can initialize the scales of Gaussian primitives in this area as a large number. Low-order geometric quantities, such as tangent and normal vectors, have already played key roles in recent works, and we further extend their application scopes in our framework. For example, to reduce floating artifacts, we truncate the gradient update of a Gaussian primitive in terms of its normal direction.
Secondly, we present an estimation method that can provide noise-robust geometric information for our framework. Specifically, we assume that each Gaussian primitive resides on an underlying surface that is locally a manifold (Lee, 2018), and we derive the analytical form of the shape operator, which contains the curvature information. We also investigate another approach, based on the notion of varifold from geometric measure theory (Allard, 1972; Simon, 1983; Menne, 2017), which in some cases outperforms our manifold-based method. Both methods are highly efficient, and thus can provide our framework with dynamic geometric priors during training.
Extensive experiments have been conducted on multiple benchmark datasets in novel view synthesis. The results demonstrate that our framework, GeoSplat, significantly improves the performance of Gaussian splatting and outperforms previous baselines.
@@ 2.1 Gaussian Splatting
The core of Gaussian splatting is to represent a 3D scene as a number of Gaussian primitives $\{\mathcal{G}^{i}_{\mathrm{obj}}\}_{i\in\mathcal{I}}$, with each primitive parameterized by its opacity $\alpha^{i}_{\mathrm{obj}}\in\mathbb{R}$, RGB color $\mathbf{c}^{i}_{\mathrm{obj}}\in\mathbb{R}^{3}$, mean vector $\bm{\mu}^{i}_{\mathrm{obj}}\in\mathbb{R}^{3}$, and covariance matrix $\bm{\Sigma}^{i}_{\mathrm{obj}}\in\mathbb{R}^{3\times 3}$. The matrix $\bm{\Sigma}^{i}_{\mathrm{obj}}$ is semipositive definite and can be decomposed as $\mathbf{R}^{i}_{\mathrm{obj}}\mathbf{S}^{i}_{\mathrm{obj}}(\mathbf{R}^{i}_{\mathrm{obj}}\mathbf{S}^{i}_{\mathrm{obj}})^{\top}$, where $\mathbf{S}^{i}_{\mathrm{obj}}=\mathrm{diag}([s^{i}_{1},s^{i}_{2},s^{i}_{3}]^{\top})$ is a diagonal scale matrix and $\mathbf{R}^{i}_{\mathrm{obj}}=[\mathbf{r}_{1}^{i};\mathbf{r}_{2}^{i};\mathbf{r}_{3}^{i}]\in\mathrm{SO}(3)$ is a rotation matrix.
@@
where $\mathbf{T}_{\mathrm{pose}}\in\mathrm{SE}(3)$ is the camera pose consisting of rotation matrix $\mathbf{W}_{\mathrm{pose}}$ and translation vector $\mathbf{b}_{\mathrm{pose}}$, and $\Pi$ is the 3D to 2D coordinate projection, with $\mathbf{J}$ its Jacobian matrix. Then, the blending operation renders the color of pixel $\mathbf{p}$ as
@@
where $\beta^{k}(\mathbf{p}):=\alpha^{k}_{\mathrm{obj}}\mathcal{N}(\mathbf{p};\bm{\mu}^{k}_{\mathrm{plane}},\bm{\Sigma}^{k}_{\mathrm{plane}})$ quantifies the radiance of a Gaussian primitive. Under this scheme, the optimization objective is to match rendered images with real images.
@@ 2.2 Problem Formulation
Our primary goal is to leverage diverse geometric priors to improve Gaussian splatting optimization, especially higher-order ones (e.g., curvature). Meanwhile, we also aim to estimate the priors in a noise-robust manner. These two aspects are largely neglected in recent explorations (Wang et al., 2024; Li et al., 2025) on geometric regularization.
@@
Gaussian primitives $\{\mathcal{G}^{i}_{\mathrm{obj}}\}_{i\in\mathcal{I}}$ reside in an underlying 3D scene formed by finitely many continuous manifold $\{\mathcal{M}_{j}\}_{j\in\mathcal{J}}$. Notably, each manifold $\mathcal{M}_{j}$ is smooth up to a zero-measure set of singular points.
@@
Compared with vanilla manifold, this new setting can accommodate typical 3D scenes, where there are multiple independent spatial objects that might not be fully smooth. The zero-measure condition is also important, permitting almost every point $\mathbf{q}\in\mathbb{R}^{3}$ in the scene to locate at a local region that can be treated as a smooth manifold. For novel view synthesis, the surface $\mathcal{M}_{j}$ is normally 2D (we denote its dimension as $D_{\mathcal{M}_{j}}=2$), embedded in the 3D ambient space $\mathcal{E}=\mathbb{R}^{3}$ ($D_{\mathcal{E}}=3)$.
@@
Most geometric quantities are locally defined, and thus fit our setting. Specifically, every Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}},i\in\mathcal{I}$ will lie on a certain underlying object surface $\mathcal{M}_{j},j\in\mathcal{J}$ at position $\mathbf{q}=\bm{\mu}^{i}_{\mathrm{obj}}$, with tangent space $\mathcal{T}_{\mathbf{q}}\mathcal{M}_{j}$ spanned by orthonormal basis $\{\mathbf{u}^{i}_{d}\}_{1\leq d\leq D_{\mathcal{M}_{j}}}$ and normal space $\mathcal{N}_{\mathbf{q}}\mathcal{M}_{j}$ formed by another basis $\{\widetilde{\mathbf{u}}^{i}_{d}\}_{D_{\mathcal{M}_{j}}+1\leq d\leq D_{\mathcal{E}}}$. These vector bases represent first-order geometric quantities, and we also aim to consider higher-order ones, e.g. principal curvatures and directions $\{(\tau_{d}^{i},\widetilde{\mathbf{w}}^{i}_{d})\}_{1\leq d\leq D_{\mathcal{M}_{j}}}$, with each pair indicating that the underlying surface is bent with a magnitude of $\tau_{d}^{i}$ along direction $\widetilde{\mathbf{w}}^{i}_{d}$. For notational convenience, we will omit the subscript $j$ of local manifold $M_{j}$, since it is unique for each primitive $\mathcal{G}^{i}_{\mathrm{obj}}$.
Another geometric tool: varifolds.
@@
We also explore adopting another tool from geometric measure theory, the notion of varifold (Allard, 1972; Simon, 1983; Menne, 2017), which is free from the smoothness assumption of manifolds. Generally speaking, a varifold is a nonnegative Radon measure (Rudin, 1987) on the product space $\mathcal{E}\times\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$ where the second space denotes the Grassmannian of $D_{\mathsf{M}}$-dimensional vector subspaces of $\mathcal{E}$ (Lee, 2018). Two classes of varifolds useful in our context are rectifiable varifolds and point cloud varifolds. A rectifiable varifold has the form $\mathcal{W}=\theta\mathcal{H}^{D_{\mathsf{M}}}_{|\mathsf{M}}\otimes\delta_{\mathcal{T}_{x}\mathsf{M}}$ where $\mathcal{H}^{D_{\mathsf{M}}}_{|\mathsf{M}}$ is the $D_{\mathsf{M}}$-dimensional Hausdorff measure restricted to a countably rectifiable set $\mathsf{M}$ (Federer, 1969; Simon, 1983), $\delta_{\mathcal{T}_{x}\mathsf{M}}$ is the Dirac measure in $\mathsf{G}_{D_{\mathsf{M}},D_{\mathcal{E}}}$ supported on the tangent plane at $x\in\mathsf{M}$, and $\theta:\mathsf{M}\rightarrow\mathbb{R}$ is a nonnegative multiplicity function with $\theta>0$ $\mathcal{H}^{D_{\mathsf{M}}}$-a.e. on $\mathsf{M}$. A point cloud $d$-varifold has the form $\mathcal{W}=\sum_{1\leq j\leq N}m_{j}\delta_{x_{j}}\otimes\delta_{P_{j}}$ where $\{x_{j}\}_{j=1}^{N}$ is a finite set of points associated with $d$-planes $\{P_{j}\}_{j=1}^{N}$ and nonnegative masses $\{m_{j}\}_{j=1}^{N}$.
@@
The notion of varifold is so flexible that it can be associated with the discrete locations of Gaussian primitives $\{\mathcal{G}^{i}_{\mathrm{obj}}\}_{i\in\mathcal{I}}$ without supposing there is any smooth interpolation. More importantly, geometric quantities valid for manifolds can be weakly or approximately defined for varifolds, such as the approximate tangent space or the (approximate) generalized second fundamental form and mean curvature. Due to the limited space of main text, we provide more details in Appendix A.2, with a concise review of varifold theory.
@@ 3 Method: Geometry-constrained Gaussian Splatting
In this section, we first present GeoSplat, a framework of geometry-constrained Gaussian splatting, which exploits higher-order geometric priors that are largely neglected by previous works. Then, we present noise-robust methods to estimate dynamic geometric priors. In contrast, previous methods obtained static priors through noise-sensitive or domain-specific techniques.
@@ 3.1.1 Curvature-guided Gaussian Initialization
We assume that the primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ locates at an underlying surface $\mathcal{M}$ [^1] With abuse of notation, we mostly denote the 2D surface as $\mathcal{M}$, potentially assuming it is a local manifold. However, it should be considered as a rectifiable set $\mathsf{M}$ in the context of rectifiable varifolds., so it is intuitive to warm up its initial shape $\mathbf{\Sigma}^{i}_{\mathrm{obj}}$ in terms of geometric priors. To begin with, we set one rotation direction as the normal vector to $\mathcal{M}$, with the corresponding scale specified as a tiny value:
$\mathbf{r}_{3}^{i}=\widetilde{\mathbf{u}}^{i}_{3},\ \ \ s^{i}_{3}=\xi_{\mathrm{min}}\ll 1.$ (3)
@@
If MAC is quite small, i.e. $\widebar{\tau}^{i}<\xi_{\mathrm{min}}$, which means the curvature is basically negligible, then we can identify that the corresponding Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ resides in an almost flat region. In this case, we upsample its position through midpoint interpolation:
@@ 3.1.2 Shape-constrained Optimization
where $\omega$ is the learning rate, $\nabla^{i}\mathfrak{L}$ denotes the gradient of loss function $\mathfrak{L}$ regarding position $\bm{\mu}^{i}_{\mathrm{obj}}$, $(\cdot)^{\top}$ denotes the orthogonal projection onto the tangent plane $\mathcal{T}_{\mathbf{q}}\mathcal{M}$, and $(\cdot)^{\perp}$ represents the projection onto the normal direction $\widetilde{\mathbf{u}}^{i}_{3}$. With this truncation, a Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ that lies on the underlying surface $\mathcal{M}$ will still stay close to it after an aggressive gradient update $\nabla^{i}\mathfrak{L}$.
@@
Another type of artifact that might degrade the performance of Gaussian splatting is the Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ with a needle-like shape (Hyung et al., 2024): covariance matrix $\mathbf{\Sigma}^{i}_{\mathrm{obj}}$ degenerates to be rank one. From a geometric perspective, the main scales $s^{i}_{1},s^{i}_{2}$ of the primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ must be regularized to ensure a rank two covariance matrix $\mathbf{\Sigma}^{i}_{\mathrm{obj}}$, so that it will have a sufficient coverage over the underlying manifold $\mathcal{M}$.
@@
with the last term to ensure the rank of matrix $\mathbf{\Sigma}^{i}_{\mathrm{obj}}$ is smaller than $3$, hence the Gaussian primitive $\mathcal{G}^{i}_{\mathrm{obj}}$ is closely attached to the surface $\mathcal{M}$. For the hinge term, we again regard the curvature information $\{\tau_{d}^{1},\tau_{d}^{2}\}$ as natural guidance for shaping the scale matrix $\mathbf{S}^{i}_{\mathrm{obj}}$. In a similar spirit, we introduce another regularization for the rotation matrix $\mathbf{R}^{i}_{\mathrm{obj}}$ as